OpenStudy (anonymous):

Solve |2y - 2| = y. A. {2/3, 2} B. {2, 3} C. {2, 5} D. {1/3, 2}

myininaya (myininaya):

Solve the following two equations 2y-2=y or 2y-2=-y

myininaya (myininaya):

Please don't just give answers. See the new conduct code thanks

myininaya (myininaya):

http://openstudy.com/code-of-conduct

myininaya (myininaya):

Are you having trouble solving 2y-2=y?